# K3 surfaces with real or complex multiplication

### v2

@@ Abstract.
Let $E$ be a totally real number field of degree $d$ and let $m\geqslant 3$ be an integer. We show that if $md\leqslant 21$ then there exists an $(m-2)$-dimensional family of complex projective $K3$ surfaces with real multiplication by $E$. Analogous results are proved for CM number fields and also for all known higher-dimensional hyperkähler manifolds.
@@ 0. Introduction
Several partial results are in the literature. Taelman [T 16] proved that if $E$ is a CM field of degree $\leqslant 20$, then there exists a complex projective $K3$ surface $X$ such that $A_{X}\simeq E$ and ${\rm dim}_{E}(T_{X,{\mathbf{Q}}})=1$. For totally real fields, several results were proved in [vG 08], Elsenhans-Jahnel [EJ 14], [EJ 16], [EJ 23] and recently in [vGS 23].
@@
Theorem A. Let $E$ be a totally real number field of degree $d$ and let $m$ be an integer with $m\geqslant 3$ and $md\leqslant 21$. Then there exists an $(m-2)$-dimensional family of complex projective $K3$ surfaces such that a very general member $X$ has the properties $A_{X}\simeq E$ and ${\rm dim}_{E}(T_{X,{\mathbf{Q}}})=m$.
@@
If $m\geqslant 2$, then there exists an $(m-1)$-dimensional family of complex projective $K3$ surfaces such that a very general member $X$ has the properties $A_{X}\simeq E$ and ${\rm dim}_{E}(T_{X,{\mathbf{Q}}})=m$.
@@
The term complex multiplication is usually reserved for the case when $A_{X}$ is a CM field and ${\rm dim}(T_{X,{\mathbf{Q}}})=1$, since this is equivalent to the Mumford-Tate group of $X$ being abelian, see [Z 83]. However, several authors use the term complex multiplication for surfaces $X$ such that $A_{X}$ is CM, also when ${\rm dim}(T_{X,{\mathbf{Q}}})>1$.
We obtain analogous results for projective hyperkähler (HK) manifolds, these have Hodge structures that are similar to the one of a K3 surface, to the extent that all the totally real fields and CM fields which are a priori possible for degree reasons actually occur (see Theorems 14.1 and 14.2 in Section 14).
To explain the strategy of the proofs, we observe that having RM or CM depends only on the rational, polarized Hodge structure $T_{X,{\mathbf{Q}}}$. Let $E$ be the algebra of Hodge endomorphisms of this Hodge structure. The ${\mathbf{Q}}$-vector space $T_{X,{\mathbf{Q}}}$ is then a module over the field $E$. Zarhin showed that this field is either totally real or a CM field, and that the quadratic form $q$ which defines the polarization has the (adjoint) property
$q(\alpha x,y)\,=\,q(x,\overline{\alpha}y)$
for all $x,y\in T_{X,{\mathbf{Q}}}$ and all $\alpha\in E$ (see Section 11.1). Here $\alpha\to\overline{\alpha}$ is the complex conjugation if $E$ is CM, and the identity if $E$ is totally real. An $E$-vector space $U$ with quadratic form $q$ has this property if and only if there exists of a hermitian (if $E$ is CM) or quadratic (if $E$ is totally real) form $h:W\times W\to E$ such that for all $x,y\in W$, we have
@@
see Lemma 4.1. Here $W=U$, considered as an $E$-vector space. To distinguish between the quadratic form (over ${\mathbf{Q}}$) $U=(U,q)$ and the hermitian form $W=(W,h)$ (over $E$), we use different notations for the vector spaces, even though they are the same; we also introduce the notation
@@
Even if we are mainly interested in the case $U=T_{X,{\mathbf{Q}}}$ it is now quite natural to consider more general Hodge structures of K3 type for which analogous facts hold. In particular the quadratic form $q$ should have signature $(2,r-2)$ where $r=\dim_{\mathbf{Q}}U$. Moreover, the Hodge structure, of weight two, on $U$ should have $\dim U^{2,0}=1$ and it should not have any non-trivial Hodge substructure. In case $E$ is totally real, there is a further condition on the signature of the eigenspaces of the $E$-action on $U\otimes_{\mathbf{Q}}\mathbf{R}$, which we discuss later (see Section 11.1).
@@
Theorem ${\mathrm{A}}^{\prime}$. Let $L$ be a lattice of rank $\rho$ and signature $(1,\rho-1)$ that allows a primitive embedding into the K3 lattice. Let $d\geqslant 1$ be an odd integer, and let $m\geqslant 3$ be such that $\rho+md=22$. Let $E$ be a totally real field of degree $d$. Then there exists an $m-2$-dimensional family of complex projective $K3$ surfaces such that a very general member $X$ is such that
$\bullet$ $A_{X}\simeq E$;
$\bullet$ ${\rm Pic}(X)\simeq L$.
@@
In more detail, after covering some basics on quadratic forms in Sections 1 - 3, we discuss the transfer of quadratic and hermitian forms from number fields to ${\mathbf{Q}}$ in Section 5. Sections 6, 7 are oriented towards answering questions (ii) and (iii); combined with the fundamental facts about K3 surfaces, the results of these sections are sufficient to prove Theorems A and B in §12. Section 6 relies on the results of [B 23]. In Section 7 we mainly use the results of Sections 1, 3 and 5. However, in the case of codimension 1 (that is, $md=21$), these methods do not suffice: instead, we apply some results of Krüskemper in [Kr 92]. The methods of [Kr 92] are then further developed in the next sections of the paper.
@@
After reviewing the necessary theory of general Hodge structures of K3 type in Section 10, we consider those associated to K3 surfaces in Section 10.3 and to HK manifolds in Section 10.4. Based on this we give the proof of Theorems A and B in Section 12. In Section 13, where Theorem ${\mathrm{A}}^{\prime}$ is proved, we obtain results on lattices (and not just on the ${\mathbf{Q}}$-vector spaces they generate). The paper concludes with results, analogous to Theorems A and B, for HK manifolds in Sections 14.1, 14.2 and for Mumford–Tate groups in Section 14.3.
@@ 1. Quadratic forms
[…]
@@
For all prime numbers $p$, we have ${\rm Br}_{2}({\bf Q}_{p})\simeq{\bf Z}/2{\bf Z}$, and ${\rm Br}_{2}(\mathbf{R})\simeq{\mathbf{Z}}/2{\mathbf{Z}}$; we identify these groups with $\{0,1\}$.
@@
(1) The sign of ${\rm det}(V)$ is $(-1)^{s}$.
(2) The image of $w(V)$ in ${\rm Br}_{2}(\mathbf{R})={\mathbf{Z}}/2{\mathbf{Z}}$ is $s(s-1)/2$ mod ${\rm mod}\,2$.
(3) If ${\rm dim}(V)=1$, or if ${\rm dim}(V)=2$ and if $p$ is a prime number such that the image of ${\rm det}(V)$ in ${\mathbf{Q}}_{p}^{\times}/{\mathbf{Q}}_{p}^{\times 2}$ is $-1$, then the image of $w(V)$ in ${\rm Br}_{2}({\mathbf{Q}}_{p})$ is 00.
@@
$\bullet$ ${\rm dim}(U)<{\rm dim}(V)-2$, or
$\bullet$ ${\rm dim}(U)={\rm dim}(V)-2$ and moreover for $p=2$ and for all prime numbers $p$ dividing ${\rm det}(U)$ or ${\rm det}(V)$ we have ${\rm det}(U)\not=-{\rm det}(V)$ in ${\mathbf{Q}}_{p}^{\times}/{\mathbf{Q}}_{p}^{\times 2}$.
@@
Proof. Let $V^{\prime}$ be a quadratic form over ${\mathbf{Q}}$ such that ${\rm dim}(V^{\prime})={\rm dim}(V)-{\rm dim}(U)$, that the determinant of $V^{\prime}$ is ${\rm det}(V).{\rm det}(U)$, that the signature of $V^{\prime}$ is $(r-r^{\prime},s-s^{\prime})$ and that $w(V^{\prime})=w(V)+w(U)+({\rm det}(V),{\rm det}(U))$; this is possible by Theorem 1.1 (ii). Therefore the invariants of $V$ and of $U\oplus V^{\prime}$ coincide; applying Theorem 1.1 (i) we conclude that $V\simeq U\oplus V^{\prime}$. ∎

### v3

@@ Abstract.
Let $E$ be a totally real number field of degree $d$ and let $m\geqslant 3$ be an integer. We show that if $md\leqslant 21$ then there exists an (m2)(m-2)-dimensional family of complex projective $K3$ surfaces with real multiplication by $E$. Analogous results are proved for CM number fields and also for all known higher-dimensional hyperkähler manifolds.
Key words and phrases:
K3 surface, Hodge structure, quadratic form, real multiplication, complex multiplication, hyperkähler manifold
2010 Mathematics Subject Classification:
14J28 (primary); 14C30, 14J42 (secondary)
@@ 0. Introduction
Several partial results are in the literature. Taelman [T 16] proved that if $E$ is a CM field of degree $\leqslant 20$, then there exists a complex projective $K3$ surface $X$ such that $A_{X}\simeq E$ and ${\rm dim}_{E}(T_{X,{\mathbf{Q}}})=1$. For totally real fields, several results were proved in [vG 08], Elsenhans–Jahnel [EJ 14], [EJ 16], [EJ 23] and recently in [vGS 25].
@@
THEOREM A. Let $E$ be a totally real number field of degree $d$ and let $m$ be an integer with $m\geqslant 3$ and $md\leqslant 21$. Then there exists an (m2)(m-2)-dimensional family of complex projective $K3$ surfaces such that a very general member $X$ has the properties $A_{X}\simeq E$ and ${\rm dim}_{E}(T_{X,{\mathbf{Q}}})=m$.
@@
If $m\geqslant 2$, then there exists an (m1)(m-1)-dimensional family of complex projective $K3$ surfaces such that a very general member $X$ has the properties $A_{X}\simeq E$ and ${\rm dim}_{E}(T_{X,{\mathbf{Q}}})=m$.
@@
The term complex multiplication is usually reserved for the case when $A_{X}$ is a CM field and ${\rm dim}_{A_{X}}(T_{X,{\mathbf{Q}}})=1$, since this is equivalent to the Mumford-Tate group of $X$ being abelian, see [Z 83]. However, several authors use the term complex multiplication for surfaces $X$ such that $A_{X}$ is CM, also when ${\rm dim}_{A_{X}}(T_{X,{\mathbf{Q}}})>1$.
We obtain analogous results for projective hyperkähler (HK) manifolds, these have Hodge structures that are similar to the one of a K3 surface, to the extent that all the totally real fields and CM fields which are a priori possible for degree reasons actually occur (see Theorems 15.1 and 15.2 in Section 15).
Before we comment on the content of the paper in more detail, let us motivate it by briefly highlighting implications of RM or CM. In a nutshell, the most striking consequence is the following from [vG 08, §7]:
If a complex manifold $X$ of dimension $2n$ has a Hodge substructure $V\subset H^{2k}(X,{\mathbf{Z}})$ of K3-type with RM or CM by a field $K$, then any non-zero $x\in K$ induces an isomorphism $\phi_{x}:V\rightarrow V$; moreover, this isomorphism gives an integral Hodge class in $H^{2n}(X\times X,{\mathbf{Z}})$ (cf. [vG 00, §2.6] or [vGS 25, §4.8]). The Hodge conjecture asserts that this class is the class of an algebraic cycle on $X\times X$, but this is not known even for the case that $X$ is a K3 surface. (In fact, in the K3 case, the Hodge conjecture holds on $X\times X$ if and only if all RM and CM Hodge classes as above are given by algebraic cycles, compare [C 14, Remark 7].) These cycles can sometimes be obtained explicitly from the graphs of certain rational self-maps (which may be of independent interest, cf. [vGS 25, §6,7]) or from covering structures as in [vGS 25, §4,5].
On the arithmetic side, there is a consequence of RM or CM relying on the Tate conjecture: Assume that the K3 surface $X$ is defined over some number field $L$; upon increasing $L$, if necessary, the Galois action of Gal(L¯/L)($\bar${L}/L) should commute with the action of the RM or CM field $K$ on the étale cohomology of $X\otimes_{L}\bar{L}$. This has strong implications on the associated $L$-series and zeta functions and can also be used to detect (or exclude) RM or CM, cf. [EJ 14], [EJ 16].
Strategy of proofs
To explain the strategy of the proofs, we observe that having RM or CM depends only on the rational, polarized Hodge structure $T_{X,{\mathbf{Q}}}$. Let $E$ be the algebra of Hodge endomorphisms of this Hodge structure. Then $E$ is a (commutative) field, and the $\mathbf{Q}$}-vector space $T_{X,{\mathbf{Q}}}$ has a structure of an $E$-vector space. Zarhin showed that the field $E$ is either totally real or a CM field, and that the quadratic form $q$ which defines the polarization has the (adjoint) property
(0.1) $q(\alpha x,y)\,=\,q(x,\overline{\alpha}y)$
for all $x,y\in T_{X,{\mathbf{Q}}}$ and all $\alpha\in E$ (see Section 12.1). Here $\alpha\mapsto\overline{\alpha}$ is the complex conjugation if $E$ is CM, and the identity if $E$ is totally real. A $\mathbf{Q}$}-vector space $U$ with quadratic form $q$ has this property if and only if there exists of a hermitian (if $E$ is CM) or quadratic (if $E$ is totally real) form $h:W\times W\to E$ such that for all $x,y\in W$, we have
@@
see Lemma 5.1. Here $W=U$ as a $\mathbf{Q}$}-vector space, but we consider $W$ as an $E$-vector space. To distinguish between the quadratic form (over $\mathbf{Q}$}) $U=(U,q)$ and the hermitian form $W=(W,h)$ (over $E$), we use different notations for the vector spaces, even though they are the same; we also introduce the notation
@@
Even if we are mainly interested in the case $U=T_{X,{\mathbf{Q}}}$ it is now quite natural to consider more general Hodge structures of K3 type for which analogous facts hold. In particular the quadratic form $q$ should have signature (2,r2)(2,r-2) where $r=\dim_{\mathbf{Q}}U$. Moreover, the Hodge structure, of weight two, on $U$ should have $\dim U^{2,0}=1$ and it should not have any non-trivial Hodge substructure. In case $E$ is totally real, there is a further condition on the signature of the eigenspaces of the $E$-action on $U\otimes_{\mathbf{Q}}\mathbf{R}$, which we discuss later (see Section 12.1).
@@
THEOREM $\mathrm{A}$$^{\prime}$. Let $L$ be a lattice of rank $\rho$ and signature $(1,\rho-1)$ that allows a primitive embedding into the K3 lattice. Let $d\geqslant 1$ be an odd integer, and let $m\geqslant 3$ be such that $\rho+md=22$. Let $E$ be a totally real field of degree $d$. Then there exists an m2m-2-dimensional family of complex projective $K3$ surfaces such that a very general member $X$ is such that
  • $\bullet$
    $A_{X}\simeq E$;
  • $\bullet$
    ${\rm Pic}(X)\simeq L$.
@@
Outline of the paper
In more detail, after covering some basics on Brauer groups and quadratic forms in Sections 1 - 4, we discuss the transfer of quadratic and hermitian forms from number fields to $\mathbf{Q}$} in Section 6. Sections 7, 8 are oriented towards answering questions (ii) and (iii); combined with the fundamental facts about K3 surfaces, the results of these sections are sufficient to prove Theorems A and B in §13. Section 7 relies on the results of [B 24a]. In Section 8 we mainly use the results of Sections 2, 4 and 6. However, in the case of codimension 1 (that is, $md=21$), these methods do not suffice: instead, we apply some results of Krüskemper in [Kr 92]. The methods of [Kr 92] are then further developed in the next sections of the paper.
@@
After reviewing the necessary theory of general Hodge structures of K3 type in Section 11, we consider those associated to K3 surfaces in Section 11.3 and to HK manifolds in Section 11.4. Based on this we give the proof of Theorems A and B in Section 13. In Section 14, where Theorem $\mathrm{A}$$^{\prime}$ is proved, we obtain results on lattices (and not just on the $\mathbf{Q}$}-vector spaces they generate). The paper concludes with results, analogous to Theorems A and B, for HK manifolds in Sections 15.1, 15.2 and for Mumford–Tate groups in Section 15.3.
1. Brauer groups
The aim of this section is to give some reminders on Brauer groups that will be useful throughout the paper. Let $k$ be a field of characteristic not 2. The Bauer group of $k$, denoted by ${\rm Br}(k)$, can be defined using central simple algebras or Galois cohomology (see for instance [GS 06], [L 05] or [S 79]). We use the additive notation for this group. In the sequel, we only need the subgroup ${\rm Br}_{2}(k)$ consisting of $x\in{\rm Br}(k)$ such that $x+x=0$. Typical elements of this group are given by (classes of) quaternion algebras.
Definition 1.1.
Let $a,b\in k^{\times}$. The quaternion algebra $(a,b)$ is the associative $k$-algebra with basis $\{1,i,j,k\}$ where $i^{2}=a,j^{2}=b$ and ij=k=jiij=k=-ji.
Example 1.2.
Over $\mathbf{R}$, the only quaternion algebras are the classical Hamilton quaternions (which form a division algebra) and the real algebra of $2\times 2$ matrices (which is also called split).
Over $\mathbf{C}$, there is only one quaternion algebra up to isomorphism, namely the complex algebra of $2\times 2$ matrices.
We only need to consider the case where $k$ is a number field or a local field. For these fields, the group ${\rm Br}_{2}(k)$ is generated by quaternion algebras. Their arithmetic is well understood in terms of local and global class field theory, see for instance the presentation in [S 79] or [V 21]. We summarize here the properties that are needed in the paper.
If $k$ is a local field, then ${\rm Br}_{2}(k)$ has two elements, and we identify it with ${\mathbf{Z}}/2{\mathbf{Z}}$. If $k$ is a number field, then the Brauer–Hasse–Noether theorem gives us the following exact sequence
$0\to\mathrm{Br}_{2}(k)\to\bigoplus_{v}\mathrm{Br}_{2}(k_{v})\to{\mathbf{Z}}/2{\mathbf{Z}}\to 0$
where the sum runs over all places $v$ of $k$. In other words, the global elements are determined by their localizations, and have to satisfy a reciprocity condition.
Another way of looking at the quaternion algebras $(a,b)$ is in terms of the associated Hilbert symbols. These are also denoted by $(a,b)$ and have the following three properties : $(a,b)$ is symmetric, i.e. $(a,b)=(b,a)$, it is bilinear (in particular, $(a,b+b^{\prime})=(a,b)+(a,b^{\prime})$), and it satisfies the relation (a,a)=0(a,-a)=0 for all $a\in k^{\times}$. See for instance [Sch 85, Corollary 11.13] for a list of useful properties, including these ones.
For $k={\bf Q}_{p}$, the field of $p$-adic numbers, the symbols $(a,b)$ can be computed using [S 77, Theorem 1, p. 39]. The following examples will be useful in the sequel.
Example 1.3.
(1) Let $a\in k^{\times}$. There is the relation (a,a)=(1,1)(a,a)=(-1,-1), and this is equivalent to (1,a)=0(-1,-a)=0. Indeed, we have (a,a)=(a,a)+(a,1)=(a,1)(a,a)=(a,-a)+(a,-1)=(a,-1) (because (a,a)=0(a,-a)=0). Therefore (a,a)=(1,1)(a,a)=(-1,-1) gives (a,1)=(1,1)(a,-1)=(-1,-1), or equivalently (a,1)+(1,1)=0(a,-1)+(-1,-1)=0 which is equivalent to (a,1)=0(-a,-1)=0.
(2) Let $a\in k^{\times}$. Then the relation (a,3a)=(2,6)(a,3a)=(-2,-6) can be reformulated as (3,2a)=0(-3,-2a)=0. Indeed, start with (a,3a)=(a,a)+(a,3)=(a,3)(a,3a)=(a,-a)+(a,-3)=(a,-3). We have (a,3a)=(2,6)(a,3a)=(-2,-6), hence (a,3)=(2,6)(a,-3)=(-2,-6) by the previous argument, equivalently (a,3)+(2,6)=0(a,-3)+(-2,-6)=0. But (2,6)=(2,2)+(2,3)=(2,3)(-2,-6)=(-2,2)+(-2,-3)=(-2,-3). Therefore we get (a,3)+(2,3)=0(a,-3)+(-2,-3)=0, equivalently (2a,3)=0(-2a,-3)=0.
Another useful property is phrased in terms of quadratic extensions of $k$.
Lemma 1.4.
Let $a,b\in k^{\times}$. Then $(a,b)=0$ if and only if $a$ is a norm of the quadratic extension $k(\sqrt{b})/k$.
To put this to work, we denote be the discriminant of a number field $E$ by $\Delta_{E}$, the norm of an extension $E/E_{0}$ by ${\rm N}_{E/E_{0}}$, and likewise the trace by ${\rm Tr}_{E/E_{0}}$.
Example 1.5.
Let $d$ be a square free integer, and set $E=Q(\sqrt{d})$. We have the following equivalences:
$d$ is a sum of two squares $\iff$ $d\in{\rm N}_{{\mathbf{Q}}(i)/{{\mathbf{Q}}}}({{\mathbf{Q}}(i)})$ $\iff$ (1,d)=0(-1,d)=0 $\iff$ $1\in{\rm N}_{E/{{\mathbf{Q}}}}(E) \iff$ $d\in{\rm N}_{E/{{\mathbf{Q}}}}(E)$ $\iff$ $\Delta_{E}\in{\rm N}_{E/{{\mathbf{Q}}}}(E)$.
@@ 2. Quadratic forms
Let $k$ be a field of characteristic not 2. A quadratic form over $k$ is by definition a pair $V=(V,q)$, where $V$ is a finite dimensional $k$-vector space and $q:V\times V\to k$ is a non-degenerate symmetric bilinear form. Every quadratic form $V$ can be diagonalized: there exist $a_{1},\dots,a_{n}\in k^{\times}$ such that $V$ is isomorphic to the diagonal quadratic form $\langle a_{1},\ldots,a_{n}\rangle$. The determinant of $V$ is by definition ${\rm det(}V)=\prod_{i}a_{i}$ in $k^{\times}/k^{\times 2}$. The Hasse invariant of $V$ is $w(V)=\sum_{i<j}(a_{i},a_{j})$ in ${\rm Br}_{2}(k)$, where $(a_{i},a_{j})$ is the class of the quaternion algebra determined by $a_{i}$ and $a_{j}$. If $W$ is another quadratic form, then
@@
Recall that, for all prime numbers $p$, we have ${\rm Br}_{2}({\bf Q}_{p})\simeq{\bf Z}/2{\bf Z}$ while also ${\rm Br}_{2}(\mathbf{R})\simeq{\mathbf{Z}}/2{\mathbf{Z}}$; we identify these groups with $\{0,1\}$.
@@
  1. (1)
    The sign of ${\rm det}(V)$ is $(-1)^{s}$.
  2. (2)
    The image of $w(V)$ in ${\rm Br}_{2}(\mathbf{R})={\mathbf{Z}}/2{\mathbf{Z}}$ is s(s1)/2s(s-1)/2 ${\rm mod}\,2$.
  3. (3)
    If ${\rm dim}(V)=1$, or if ${\rm dim}(V)=2$ and if $p$ is a prime number such that the image of ${\rm det}(V)$ in ${\mathbf{Q}}_{p}^{\times}/{\mathbf{Q}}_{p}^{\times 2}$ is 1-1, then the image of $w(V)$ in ${\rm Br}_{2}({\mathbf{Q}}_{p})$ is 0.
@@
Remark 2.2.
In particular, in dimension 1 and 2, Theorem 2.1 implies the following classification of quadratic forms over $\mathbf{Q}$} by means of the following invariants:
  1. (1)
    In rank $1$, the determinant suffices (and the signature, but the signature is the sign of the determinant here);
  2. (2)
    In rank $2$, determinant and Hasse invariant suffice (as the signature is determined by the Hasse invariant over the real numbers).
@@
  • $\bullet$
    ${\rm dim}(U)<{\rm dim}(V)-2$, or
  • $\bullet$
    ${\rm dim}(U)={\rm dim}(V)-2$ and moreover for $p=2$ and for all prime numbers $p$ dividing ${\rm det}(U)$ or ${\rm det}(V)$ we have ${\rm det}(U)\not=-{\rm det}(V)$ in ${\mathbf{Q}}_{p}^{\times}/{\mathbf{Q}}_{p}^{\times 2}$.
@@
Proof. Let $V^{\prime}$ be a quadratic form over $\mathbf{Q}$} such that ${\rm dim}(V^{\prime})={\rm dim}(V)-{\rm dim}(U)$, that the determinant of $V^{\prime}$ is ${\rm det}(V)\cdot{\rm det}(U)$, that the signature of $V^{\prime}$ is $(r-r^{\prime},s-s^{\prime})$ and that $w(V^{\prime})=w(V)+w(U)+({\rm det}(V),{\rm det}(U))$; this is possible by Theorem 2.1 (ii). Therefore the invariants of $V$ and of $U\oplus V^{\prime}$ coincide; applying Theorem 2.1 (i) we conclude that $V\simeq U\oplus V^{\prime}$. ∎